# Learning Collective Variables for Enhanced Sampling from BioEmu with Time-Lagged Generation

### v2

@@ Abstract
Molecular dynamics is crucial for understanding molecular systems but its applicability is often limited by the vast timescales of rare events like protein folding. Enhanced sampling techniques overcome this by accelerating the simulation along key reaction pathways, which are defined by collective variables (CVs). However, identifying effective CVs that capture the slow, macroscopic dynamics of a system remains a major bottleneck. This work proposes a novel framework coined BioEmu-CV that learns these essential CVs automatically from BioEmu, a recently proposed foundation model for generating protein equilibrium samples. In particular, we re-purpose BioEmu to learn time-lagged generation conditioned on the learned CV, i.e., predict the distribution of molecular states after a certain amount of time. This training process promotes the CV to encode only the slow, long-term information while disregarding fast, random fluctuations. We validate our learned CV on fast-folding proteins with two key applications: (1) estimating free energy differences using on-the-fly probability enhanced sampling and (2) sampling transition paths with steered molecular dynamics. Our empirical study also serves as a new systematic and comprehensive benchmark for MLCVs on fast-folding proteins larger than Alanine Dipeptide.
@@ 1 Introduction
Many problems in biomolecules reduce to understanding the complex behavior of how molecules move and change conformations, such as estimating the drug-target binding affinities (de2016role; abel2017advancing) or the folding time of proteins (piana2012protein; spotte2022toward). Molecular dynamics (MD) simulation is one of the principal ways to study the movements in molecular systems, describing the microscopic movements by integrating differential equations with femtosecond ($10^{-15}$) time step. However, events like protein folding often happen in a much larger time-scale, from microseconds ($10^{-6}$) to milliseconds ($10^{-3}$). Due to this timescale gap, observing desired events in naive MD requires integrating over computationally infeasible number of step, which is considered unrealistic for most real world problems.
To overcome this timescale limitation, various enhanced sampling techniques have been studied (henin2022enhanced). Metadynamics (barducci2011metadynamics) applies biasing force to the molecular systems to encourage transitions, while replica-exchange MD (sugita1999replica) exchanges configurations between parallel simulations at different temperatures. Also, accelerated MD (hamelberg2004accelerated) globally boosts the potential energy surface to overcome energy barriers. Their core idea is to add biasing forces throughout the simulation to guide exploration and transitions in the molecular state space, without breaking the equilibrium conditions. Importantly, this biasing forces are computed based on a low-dimensional representation, denoted as the collective variables (CVs). Therefore, CVs well encoding the slow degree of freedom will add biases to numerously seen CVs, resulting exploration and transition to unseen and less visited states.
However, CVs have been mainly hand-crafted by domain expertise, e.g., selection of specific backbone dihedral angles for Alenine Dipeptide, which may miss slow modes and has been limited to small systems. To resolve this issue, several works have considered machine-learned CVs (MLCVs) from MD trajectory data. Supervised methods use ground-truth state definitions and pseudo-labels (bonati2020data; trizio2021enhanced), while self-supervised methods rely on time-lagged data to encode the dynamics information (bonati2021deep; wehmeyer2018time; hernandez2018variational). While prior works have shown promising results on small systems such as alanine dipeptide, they currently lack the ability to scale to larger systems such as proteins and lack standardized, side-by-side systematic comparison.
@@
  • We propose a framework extracting collective variables from the time-lagged condition representations of a frozen foundation model.
  • We extensively benchmark MLCVs with two downstream tasks for the slow degree of freedom: free energy difference estimation and transition path sampling for proteins in explicit water solvent simulations. Furthermore, we demonstrate that our BioEmu-CV shows competitive performance in these tasks as well as meta-stable state discrimination.
@@ 3 Method
Table 1: Quantitative results of 1 $\bm{\mu}$s OPES simulations for three fast-folding proteins in explicit water solvent. We report the reference free energy difference ($\Delta F_{\text{ref}}$) from the full dataset, the average free energy difference ($\Delta F$) from simulations, the absolute error between two energy values ($|\Delta F_{\text{ref}}-\Delta F|$), and the potential of mean force (PMF) MAE. Results are averaged over four simulations for Chignolin, and three simulations for Trp-cage and BBA. We mark not applicable (N/A) for CVs that fail at meta-stable state discrimination or show a different sign with the reference value.
@@
Figure 2: Free energy (top) and PMF (bottom) estimation from 1 $\mu s$ OPES simulations for Chignolin. We average over four seeds. Green dotted lines indicate the reference value, and blue lines refer to free energy difference during the OPES simulations. Solid lines refer to the mean, and shaded areas are the standard deviation. DeepTICA shows negative values beyond -1 in PMF, even normalized to the DESRES trajectory, falls short of accurately capturing the reference PMF.
@@ 4 Experiment results
Evaluation setup. We consider three fast-folding protein in explicit water solvent from lindorff2011fast; Chignolin, Trp-cage, and BBA. For more detail on protein data, refer to Appendix A. We quantitatively evaluate MLCVs with two downstream tasks that require encoding the slow degree of freedom: (i) free energy difference estimation, and (ii) transition path sampling. We use on-the-fly probability enhanced simulation (invernizzi2020rethinking, OPES) and CV-steered MD simulation (izrailev1999steered; fiorin2013using, SMD) for each task, respectively. Furthermore, we extensively investigate the interpretability of CVs with respect to input descriptors with the sensitivity analysis, and analyze meta-stable state discrimination with secondary structures known to be present in the folded states following fu2024collective.
Baselines. We compare our method against self-supervised CV learning methods: DeepTICA (bonati2021deep), time-lagged autoencoder (wehmeyer2018time, TAE), and variational dynamics encoder (hernandez2018variational, VDE). Since protein types and simulation configuration vary significantly by methods and lack systematic comparison, we train from scratch on identical data and time-lag with baselines using the mlcolvar package (bonati2023unified). We use pairwise $C_{\alpha}$ distances for inputs to ensure rotation and translation invariance following prior works (trizio2021enhanced; bonati2021deep), and fix CVs dimensionality to one. All MLCVs are normalized to the range $\left[-1,1\right]$ on the full DESRES trajectory after training for simplicity and visibility, with the sign assigning the folded state to one. For more detail on baselines, refer to Appendix B.
@@ 4.1 Free energy difference estimation
Evaluation criteria. Following yang2024learning, we measure (i) the convergence of the estimated reaction free energy of folding ($\Delta F$), and (ii) the quality of the potential of mean force (PMF). We compute the two meta-stable state basins, folded and unfolded states, by dividing the CV range in half as in prior work. The reference values for each protein are computed from MLCVs of the full DESRES trajectory with the multi-state Bennett acceptance ratio analysis (shirts2008statistically, MBAR). In short, CVs well encoding the slow degree of freedom are expected to have similar distributions between the 1 $\mu s$ OPES simulation and the long naive reference $\sim$100 $\mu s$ DESRES simulation. Note that we exclude results on MLCVs that fail on the basic criterion of discriminating between the folded and unfolded states; see more detail in Section 4.4. Since inadequate CVs produce non-converged biases and misleading PMFs, OPES cannot reliably drive transition, and its PMFs are meaningless. Results are averaged over four runs for Chignolin and three runs for Trp-cage and BBA, with one outlier excluded from the original four following yang2024learning. For more details on evaluation metric and outlier criterion, refer to Appendix B.
@@
Figure 3: 3D Visualization of transition paths. The sampled folding pathways of Chignolin, Trp-cage, and BBA by steered MD with BioEmu-CV. For simplicity, we visualize the $C_{\alpha}$ coordinates.
Results. In Table 1, we report the reference free energy difference ($\Delta F_{\text{ref}}$), free energy difference averaged over simulations ($\Delta F$), and the potential mean force (PMF) MAE. In Figure 2, we additionally plot the free energy difference throughout the OPES simulations. For the full results, please refer to Appendix C. In the following, we interpret the results in both quantitative and qualitative aspects. While most MLCVs fairly converge for Chignolin and work for larger proteins, VDE fails to scale to large proteins. Additionally, TAE appears to converge for Trp-cage, but only because it samples exclusively the unfolded state. As shown in Figure 10, the folded state is barely visited resulting convergence at an incorrect $\Delta F$, with the wrong sign relative to the reference value. Similarly, DeepTICA samples the folded state with high deviation even though the outlier was removed, resulting fluctuation in the PMF. We have clarified more interpretation in detail at Appendix C.
Figure 4: Transition paths projected onto TICA coordinates, sampled by MLCV steered MD. The white star and circle each refer to the representative unfolded and folded state, with each paths colored differently. The red convex hull is the folded states defined as RMSD from the representative folded state with cut off 2, 2, and 2.5 Å for Chignolin, Trp-cage, and BBA, respectively.
@@ 4.2 Transition path sampling
Steered molecular dynamics. We sample transition paths with steered molecular dynamics (izrailev1999steered; fiorin2013using, SMD). SMD is an enhanced sampling technique that steers the protein conformation towards a pre-defined target state with a time-dependent biasing force. CVs that well encode the slow degree of freedom would produce low-energy transition paths when used with SMD. For each protein, we do 16 NVT simulations for 500 ps using the Langevin Integrator with 1 fs time step. For more detail about SMD, refer to Appendix B.
@@
Results. In Table 2, BioEmu-CV shows the best results in hitting the target meta-stable states while achieving low transition state energies. Additionally in Figure 3, we visualize the $C_{\alpha}$ atoms in the folding process under BioEmu-CV steered MD for qualitative assessment. For the visualization of baseline CV-steered MD, refer to Appendix C. In Figure 4, we also visualize the sampled transition paths of three fast-folding proteins from SMD simulations, by projecting them onto time-lagged independent component analysis (molgedey1994separation, TICA) made from the full DESRES trajectory. We find that transition paths from our MLCVs show better target reaching and follow the region with a low PMF value defined on two time-lagged independent components.
@@ 4.3 Interpretability of CVs
Figure 5: MLCV sensitivity to $C_{\alpha}$-wise distances. We plot the top $C_{\alpha}$-wise distances with the highest sensitivity for each MLCV, where the $x$ and $y$ axis each denotes the residue index for input distances and the corresponding sensitivity value. For each sensitivity plot, we visualize top sensitive distances in the unfolded and folded state, with colors highlighted in the sensitivity plot.
@@ 4.4 Meta-stable state discrimination analysis
Meta-stable state discrimination. First, we verify whether the CVs truly distinguish the folded and unfolded states of proteins. In Figure 6, we color each TICA coordinates with its MLCVs. We use a time lag of $\tau=1000$ for BBA and $\tau=10$ otherwise for TICA plots. While most methods distinguish the folded and unfolded states, prior works tend to fail at BBA. In Table 3, we report the average MLCVs for the folded and unfolded states gathered from the full DESRES trajectory. Noticeably, VDE shows similar values for Trp-cage and BBA. We plot the detailed distribution of MLCVs for the folded and unfolded states in Appendix C, where VDE fails at meta-stable state discrimination showing almost identical results between MLCVs of the folded and unfolded states.
@@
Chignolin analysis. We also analyze CVs for Chignolin with well-known descriptors, the committor function. The committor function provides a quantitative measure of progress along the folding pathway on a scale from $[0,1]$. We use the committor function from kang2024computing, estimated in a data-driven manner. In Table 4, we report the Pearson correlation between the committor function and MLCVs. DeepTICA shows relatively weak correlation with the committor values, whereas other methods demonstrate stronger alignment. For more detail on Chignolin descriptors, refer to Appendix B.
@@
Secondary structure analysis. Finally, we analyze the correlation between CVs and the secondary structure of proteins. As seen in Table 5, secondary structure are mostly observed in the folded state clearly distinguished from the unfolded states (lindorff2011fast), therefore CVs should capture their presence to describe the folding dynamics (ahalawat2018assessment). We compute the residue-level secondary structure using the dictionary of secondary structures in proteins (kabsch1983dictionary; mcgibbon2015mdtraj, DSSP), which classifies structures based on hydrogen-bond patterns. In Figure 7, we plot the MLCV distribution against the known secondary structures in BBA. While the folded states of BBA contain one $\alpha$-helix and two $\beta$-sheets (lindorff2011fast), TAE and VDE encode nearly identical CVs regardless of the secondary structures in contrast to DeepTICA and BioEmu-CV successfully separating them.
@@
Figure 7: Secondary structure analysis on BBA. (Top) Folded-state visualization of the BBA protein, with $\alpha$-helix and $\beta$-sheet highlighted in orange and green, respectively. We also plot the MLCVs distribution for the residues included in the secondary structures, for residues showing specific secondary structures in the folded state. C, H, and E each correspond to coil (irregular elements), $\alpha$-helix, and $\beta$-sheet secondary structures. TAE and VDE fails to assigns distinct values to helix and sheet structures. (Bottom) Visualizations and MLCVs on states in the DESRES trajectory having partial secondary structures. Residues are colored according to the DSSP value. DeepTICA and our MLCVs align with the secondary structures.

### v3

# Learning Collective Variables from BioEmu with Time-Lagged Generation

@@ Abstract
Molecular dynamics is crucial for understanding molecular systems, but its applicability is often limited by the vast timescales of rare events like protein folding. Enhanced sampling techniques overcome this by accelerating the simulation along key reaction pathways, which are defined by collective variables (CVs). However, identifying effective CVs that capture the slow, macroscopic dynamics of a system remains a major bottleneck. This work proposes a novel framework coined BioEmu-CV that learns these essential CVs automatically from BioEmu, a recently proposed foundation model for generating protein equilibrium samples. In particular, we re-purpose BioEmu to learn time-lagged generation conditioned on the learned CV, i.e., predict the distribution of molecular states after a certain amount of time. This training process promotes the CV to encode only the slow, long-term information while disregarding fast, random fluctuations. We validate our learned CV on fast-folding proteins with two key applications: (1) estimating free energy differences using on-the-fly probability enhanced sampling and (2) sampling transition paths with steered molecular dynamics. Our empirical study also serves as a new systematic and comprehensive benchmark for MLCVs on fast-folding proteins larger than Alanine Dipeptide.
@@ 1 Introduction
22footnotetext: Corresponding Author.
Many problems in biomolecules reduce to understanding the complex dynamics of molecular motion and conformational changes, such as estimating drug-target binding affinities (de2016role; abel2017advancing) or the folding times of proteins (piana2012protein; spotte2022toward). Molecular dynamics (MD) simulation is one of the principal methods for studying molecular motion, describing microscopic dynamics by integrating differential equations with femtosecond ($10^{-15}$) time steps. However, events like protein folding often happen in a much larger timescale, from microseconds ($10^{-6}$) to milliseconds ($10^{-3}$). Due to this timescale gap, observing desired events in naive MD requires integrating over a computationally infeasible number of steps, which is considered unrealistic for most real-world problems.
To overcome this timescale limitation, various enhanced sampling techniques have been studied (henin2022enhanced). Metadynamics (barducci2011metadynamics) applies biasing force to the molecular systems to encourage transitions, while replica-exchange MD (sugita1999replica) exchanges configurations between parallel simulations at different temperatures. Also, accelerated MD (hamelberg2004accelerated) globally boosts the potential energy surface to overcome energy barriers. Their core idea is to add biasing forces throughout the simulation to guide exploration and transitions in the molecular state space, without breaking the equilibrium conditions. Importantly, these biasing forces are computed based on a low-dimensional representation, denoted as the collective variables (CVs). Therefore, CVs well encoding the slow degree of freedom will add biases to numerous seen CVs, resulting in exploration and transition to unseen and less visited states.
However, CVs have been mainly hand-crafted by domain expertise, e.g., selection of specific backbone dihedral angles for Alenine Dipeptide, which may miss slow modes and has been limited to small systems. To resolve this issue, several works have considered machine-learned CVs (MLCVs) from MD trajectory data. Supervised methods use ground-truth state definitions and pseudo-labels (bonati2020data; trizio2021enhanced), while self-supervised methods rely on time-lagged data to encode the dynamics information (bonati2021deep; wehmeyer2018time; hernandez2018variational). While prior works have shown promising results on small systems such as alanine dipeptide, they currently lack the ability to scale to larger systems such as proteins and lack a standardized, side-by-side systematic comparison.
@@
  • We propose a framework for extracting collective variables from the time-lagged condition representations of a frozen foundation model.
  • We extensively benchmark MLCVs with two downstream tasks for the slow degree of freedom: free energy difference estimation and transition path sampling for proteins in explicit water solvent simulations.
@@ 3 Method
Table 1: Quantitative results of 1 $\bm{\mu}$s OPES simulations for three fast-folding proteins in explicit water solvent. We report the reference free energy difference ($\Delta F_{\text{ref}}$) from the data, the average free energy difference ($\Delta F$) from simulations, the absolute error between two energy values ($|\Delta F_{\text{ref}}-\Delta F|$), and the potential of mean force (PMF) MAE. Results are averaged over four simulations for Chignolin, and three simulations for Trp-cage and BBA. We mark not applicable (N/A) for CVs that fail at meta-stable state discrimination or show a different sign with the reference value.
@@
Figure 2: Free energy (top) and PMF (bottom) estimation from 1 $\mu s$ OPES simulations for Chignolin. We average over four simulations. Green dotted lines indicate the reference value, and blue lines refer to the free energy difference during the OPES simulations. Solid lines refer to the mean, and shaded areas are the standard deviation. DeepTICA shows negative values beyond -1 in PMF even when normalized, and falls short of accurately capturing the reference PMF.
@@ 4 Experiment results
Evaluation setup. We consider three fast-folding protein in explicit water solvent from lindorff2011fast; Chignolin, Trp-cage, and BBA. For more details on protein data, refer to Appendix A. We quantitatively evaluate MLCVs with two downstream tasks that require encoding the slow degree of freedom: (i) free energy difference estimation, and (ii) transition path sampling. We use on-the-fly probability enhanced simulation (invernizzi2020rethinking, OPES) and CV-steered MD simulation (izrailev1999steered; fiorin2013using, SMD) for each task, respectively. Furthermore, we extensively investigate the interpretability of CVs with respect to input descriptors with the sensitivity analysis, and analyze meta-stable state discrimination with secondary structures known to be present in the folded states following fu2024collective. For ablation experiment results regarding architecture components, please refer to Appendix B.
Baselines. We compare our method against self-supervised CV learning methods: DeepTICA (bonati2021deep), time-lagged autoencoder (wehmeyer2018time, TAE), and variational dynamics encoder (hernandez2018variational, VDE). Since protein types and simulation configuration vary significantly by methods and lack systematic comparison, we train from scratch on identical data and time-lag with baselines using the mlcolvar package (bonati2023unified). We use pairwise $C_{\alpha}$ distances for inputs to ensure rotation and translation invariance following prior works (trizio2021enhanced; bonati2021deep), and fix CVs dimensionality to one. All MLCVs are normalized to the range $\left[-1,1\right]$ on the full DESRES trajectory after training for simplicity and visibility, with the sign assigning the folded state to positive. For more details on baselines, refer to Appendix C.
@@ 4.1 Free energy difference estimation
Evaluation criteria. Following yang2024learning, we measure (i) the convergence of the estimated reaction free energy of folding ($\Delta F$), and (ii) the quality of the potential of mean force (PMF). We compute the two meta-stable state basins, folded and unfolded states, by dividing the CV range in half as in prior work. The reference values for each protein are computed from MLCVs of the full DESRES trajectory with the multi-state Bennett acceptance ratio analysis (shirts2008statistically, MBAR). In short, CVs well encoding the slow degree of freedom are expected to have similar distributions between the 1 $\mu s$ OPES simulation and the long naive reference $\sim$100 $\mu s$ DESRES simulation. Note that we exclude results on MLCVs that fail on the basic criterion of discriminating between the folded and unfolded states; see more detail in Section 4.4. Since inadequate CVs produce non-converged biases and misleading PMFs, OPES cannot reliably drive transition, and its PMFs are meaningless. Results are averaged over four runs for Chignolin and three runs for Trp-cage and BBA, with one outlier excluded from the original four following yang2024learning. For more details on the simulation, evaluation, and outlier criterion, refer to Appendix C.
@@
Figure 3: 3D Visualization of transition paths. The sampled folding pathways of Chignolin, Trp-cage, and BBA by steered MD with BioEmu-CV. We visualize only the $C_{\alpha}$ for simplicity.
Results. In Table 1, we report the reference free energy difference ($\Delta F_{\text{ref}}$), free energy difference averaged over simulations ($\Delta F$), and the potential mean force (PMF) MAE. In Figure 2, we additionally plot the free energy difference throughout the OPES simulations. For the full results, please refer to Appendix D. In the following, we interpret the results in both quantitative and qualitative aspects. While most MLCVs fairly converge for Chignolin and work for larger proteins, VDE fails to scale to large proteins. Additionally, TAE appears to converge for Trp-cage, but only because it samples the unfolded state. As shown in Figure 14, the folded state is barely visited, resulting in convergence at an incorrect $\Delta F$, with the wrong sign relative to the reference value. Similarly, DeepTICA samples the folded state with high deviation even though the outlier was removed, resulting in fluctuation in the PMF. We have clarified more interpretations in detail in Appendix D.
Figure 4: Qualitative results on transition path sampling. (Left) Transition paths from MLCV-steered MD projected onto TICA coordinates. The white star and circle each refer to the representative unfolded and folded state. The red convex hulls indicate the folded state regions based on RMSD cutoffs from the representative folded state with cut off 2, 2, and 2.5 Å for Chignolin, Trp-cage, and BBA, respectively. (Right) Diverse intermediate states for Chignolin transition paths using BioEmu-CV steered MD, with white diamonds representing step 2,400 on TICA coordinates.
@@ 4.2 Transition path sampling
Steered molecular dynamics. We sample transition paths with steered molecular dynamics (izrailev1999steered; fiorin2013using, SMD). SMD is an enhanced sampling technique that steers the protein conformation towards a pre-defined target state with a time-dependent biasing force. CVs that well encode the slow degree of freedom would produce low-energy transition paths when used with SMD. For each protein, we do 16 NVT simulations for 500 ps using the Langevin Integrator with 1 fs time step. For more detail about SMD, refer to Appendix C.
@@
Results. In Table 2, BioEmu-CV shows the best results in hitting the target meta-stable states while achieving low transition state energies. Additionally in Figure 3, we visualize the $C_{\alpha}$ atoms in the folding process under BioEmu-CV steered MD for qualitative assessment. For the visualization of baseline CV-steered MD, refer to Appendix D. In Figure 4, we also visualize the sampled transition paths of three fast-folding proteins from SMD simulations, by projecting them onto time-lagged independent component analysis (molgedey1994separation, TICA) made from the full DESRES trajectory. We find that transition paths from our MLCVs show better target reaching and follow the region with a low PMF value defined on two time-lagged independent components.
@@ 4.3 Interpretability of CVs
Figure 5: MLCV sensitivity to $C_{\alpha}$-wise distances. We plot the top $C_{\alpha}$-wise distances with the highest sensitivity for each MLCV, where the $x$ and $y$ axes each denote the residue index for input distances and the corresponding sensitivity value. For each sensitivity plot, we visualize the top sensitive distances in the unfolded and folded states, with colors highlighted in the sensitivity plot.
@@ 4.4 Meta-stable state discrimination analysis
Meta-stable state discrimination. First, we verify whether the CVs truly distinguish the folded and unfolded states of proteins. In Figure 6, we color each TICA coordinates with its MLCVs. We use a time lag of $\tau=1000$ for BBA and $\tau=10$ otherwise for TICA plots. While most methods distinguish the folded and unfolded states, prior works tend to fail at BBA. In Table 3, we report the average MLCVs for the folded and unfolded states gathered from the full DESRES trajectory. Noticeably, VDE shows similar values for Trp-cage and BBA. We plot the detailed distribution of MLCVs for the folded and unfolded states in Appendix D, where VDE fails at meta-stable state discrimination showing almost identical results between MLCVs of the folded and unfolded states.
@@
Chignolin analysis. We also analyze CVs for Chignolin with well-known descriptors, the committor function. The committor function provides a quantitative measure of progress along the folding pathway on a scale from $[0,1]$. We use the committor function from kang2024computing, estimated in a data-driven manner. In Table 4, we report the Pearson correlation between the committor function and MLCVs. DeepTICA shows relatively weak correlation with the committor values, whereas other methods demonstrate stronger alignment. For more details on Chignolin descriptors, refer to Appendix C.
Secondary structure analysis. Finally, we analyze the correlation between CVs and the secondary structure of proteins. As seen in Table 5, secondary structures are mostly observed in the folded state clearly distinguished from the unfolded states (lindorff2011fast), therefore CVs should capture their presence to describe the folding dynamics (ahalawat2018assessment).
@@
We compute the residue-level secondary structure using the dictionary of secondary structures in proteins (kabsch1983dictionary; mcgibbon2015mdtraj, DSSP), which classifies structures based on hydrogen-bond patterns. In Figure 7, we plot the MLCV distribution against the known secondary structures in BBA. While the folded states of BBA contain one $\alpha$-helix and two $\beta$-sheets (lindorff2011fast), TAE and VDE encode nearly identical CVs regardless of the secondary structures in contrast to DeepTICA and BioEmu-CV successfully separating them.
@@
Figure 7: Secondary structure analysis on BBA. (Top) Folded-state visualization of the BBA protein, with $\alpha$-helix and $\beta$-sheet highlighted in orange and green, respectively. We also plot the MLCVs distribution for the residues included in the secondary structures, for residues showing specific secondary structures in the folded state. Letters C, H, and E each correspond to coil (irregular elements), $\alpha$-helix, and $\beta$-sheet secondary structures. TAE and VDE fail to assign distinct values to helix and sheet structures. (Bottom) Visualizations and MLCVs on states in the DESRES trajectory having partial secondary structures. Residues are colored according to the DSSP value. DeepTICA and BioEmu-CV align with the secondary structures.